### v1

@@ 4.4 Characterization and Recognition
Proof 4.23 (Proof Sketch).
@@
We analyze the structure of $G$ and we see that either $G$ is a graph that has $\operatorname{pqn}(G)=1$ due to a result from Section 4.2, or $G$ has some $\mathcal{F}_{i}$ for $i\in\{1,\ldots,8\}$ as a minor.
@@
If $G$ is a tree, then $\operatorname{pqn}(G)=1$ by Lemma 4.4. If $G$ contains exactly one cycle $C$, then $G-E(C)$ is a forest $F$, and we consider each component of $F$ as rooted at $C$. If every component is an isolated vertex or a rooted star, then $\operatorname{pqn}(G)=1$ by Lemma 4.10. If some component is more complex than a rooted star or a rooted caterpillar, then $G$ contains $\mathcal{F}_{4}$ as a minor. Now suppose some component $K$ is a rooted caterpillar. If all other components are isolated vertices, then $\operatorname{pqn}(G)=1$ by Lemma 4.12. If there are two further non-trivial components, then $G$ contains $\mathcal{F}_{5}$ as a minor. So assume that there is exactly one further non-trivial component $K^{\prime}\neq K$. If the roots of $K$ and $K^{\prime}$ have distance at least $3$ in one direction along $C$, then $G$ contains $\mathcal{F}_{8}$ as a minor. If none of the above applies, either $C$ is a triangle and $\operatorname{pqn}(G)=1$ by Corollary 4.16, or $C$ is a quadrangle and the roots of $K$ and $K^{\prime}$ are opposite on $C$, and $\operatorname{pqn}(G)=1$ by Lemma 4.14. The case where $G$ contains at least two cycles remains. If $G$ has two edge-disjoint cycles, then $G$ contains $\mathcal{F}_{2}$ as a minor. If two cycles have at least two vertices but no edge in common, then $G$ contains $\mathcal{F}_{7}$ as a minor. If any two cycles share at least one edge, then either $G=K_{2,3}$ or $G=K_{4}-e$ and we have $\operatorname{pqn}(G)=1$ by Lemmas 4.17 and 4.18, or $G$ contains $\mathcal{F}_{1}$, $\mathcal{F}_{6}$, or $\mathcal{F}_{7}$ as a minor.
@@
Theorem 4.22 implies that the structure of a graph $G$ with $\operatorname{pqn}(G)=1$ is quite limited. In fact, forbidden minors $\mathcal{F}_{1},\mathcal{F}_{2}$ and $\mathcal{F}_{3}$ imply that $G$ can contain at most three cycles. More precisely, if $G$ contains two cycles, it is necessarily a $K_{2,3}$ or a $K_{4}-e$ since forbidden minors $\mathcal{F}_{6}$ and $\mathcal{F}_{7}$ forbid any other edge. Moreover, if $G$ contains a single cycle, forbidden minor $\mathcal{F}_{4}$ implies that no non-trivial tree (i.e., a tree that is more than a caterpillar) can be attached to the cycle. Finally, forbidden minors $\mathcal{F}_{5}$ and $\mathcal{F}_{8}$ dictate how legs can be combined with an attached caterpillar. $\mathcal{F}_{5}$ implies that at most one vertex of the cycle can have legs whereas $\mathcal{F}_{8}$ implies that, for a cycle of length $4$, legs can only be attached to the vertex opposite of the vertex attached to the caterpillar whereas for cycles of lengths larger than $4$, no legs are allowed. These observations in conjunction with Lemmas 4.4, 4.6, 4.10, 4.14, 4.16, 4.17 and 4.18 imply the following:
@@ 5 PQ-Layouts of Graphs with Bounded Pathwidth and Treewidth
We present results for graphs of bounded pathwidth and bounded treewidth. We assume familiarity with these concepts; otherwise see Appendix C for formal definitions.
@@
To define the partition of $E$ into $p+1$ priority queues $\mathcal{P}_{1},\ldots,\mathcal{P}_{p+1}$, we consider a proper vertex coloring of $G$ with $\chi(G)=\omega(G)=p+1$ colors. This exists, as $G$ is an interval graph. Now for $c=1,\ldots,p+1$ let $\mathcal{P}_{c}$ be the set of all edges in $G$ whose right endpoint in the vertex ordering has color $c$. In fact, each $\mathcal{P}_{c}$ contains none of the forbidden configurations in Figure 2, and hence is a priority queue independent of the edge-weights, since in each forbidden configuration in Figure 2 there would be an edge in $G$ between the right endpoints $v_{r}$ and $v_{r^{\prime}}$ of the two edges $e$ and $e^{\prime}$ forming the forbidden configuration. But this would imply different colors for $v_{r}$ and $v_{r^{\prime}}$ and therefore different priority queues for $e$ and $e^{\prime}$.
We remark that Corollary 3.5 implies that there are graphs with pathwidth $p$ and priority queue number at least $\left\lfloor\frac{3-\sqrt{5}}{8}(p+1)\right\rfloor$ as $K_{p+1}$ has pathwidth $p$.
@@
We define a vertex ordering $\sigma$ of a $2$-tree rooted at $e^{*}=v^{\prime}v^{\prime\prime}$ to be left-growing if no vertex (different from $v^{\prime}$, $v^{\prime\prime}$) appears to the right of both its parents. First we show that it is enough to force a $p$-inversion (and hence $\operatorname{pqn}(G)\geq p$) in every left-growing vertex ordering.
@@
Figure 5: Illustration of the $2$-tree $H_{k}$ with edge-weighting $w_{k}$, for $k\geq 2$.
We then construct a sequence $\langle H_{1},w_{1}\rangle,\langle H_{2},w_{2}\rangle,\ldots$ of edge-weighted rooted $2$-trees, so that for every $k$, every left-growing vertex ordering of $\langle H_{k},w_{k}\rangle$ contains a $k$-inversion. The construction of $\langle H_{k},w_{k}\rangle$ is recursive, starting with $\langle H_{1},w_{1}\rangle$ being just a single edge of any weight. For $k\geq 2$, we start with $d=k^{2}$ vertices $a_{1},\ldots,a_{d}$ stacked onto the root edge $e^{*}=v^{\prime}v^{\prime\prime}$, and stacking a vertex $b^{\prime}_{i}$ onto each $v^{\prime}a_{i}$, as well as a vertex $b^{\prime\prime}_{i}$ onto $v^{\prime\prime}a_{i}$ for $i=1,\ldots,d$. All these edges have weight 00. Then, each $b^{\prime}_{i}a_{i}$ and $b^{\prime\prime}_{i}a_{i}$ is used as the base edge of two separate copies of $\langle H_{k-1},w_{k-1}\rangle$, where however the edge-weights $w_{k-1}$ of each copy are scaled and shifted to be in a specific half-open interval $[x,y)\subset\mathbb{R}$ depending on the current base edge ($b^{\prime}_{i}a_{i}$ or $b^{\prime\prime}_{i}a_{i}$). See Figure 5 for an illustration.
By induction, there is a $(k-1)$-inversion in each such copy of $H_{k-1}$. Using the pigeon-hole principle, and the Erdős–Szekeres theorem, we then identify an edge $a_{i}v^{\prime}$ (or $a_{i}v^{\prime\prime}$) which forms a $k$-inversion either together with $(k-1)$-inversion of one copy of $H_{k-1}$, or with $k-1$ edges, each from the $(k-1)$-inversion of a different copy of $H_{k-1}$.
To finish the proof, it is then enough to take $\langle H_{p},w_{p}\rangle$ as constructed above, and then apply 1 to obtain the desired edge-weighted $2$-tree $\langle G_{p},w_{p}\rangle$ with $\operatorname{pqn}(G_{p},w_{p})\geq p$.
@@
We observe that $H_{k}$ in the proof of Theorem 5.3 contains $4$ copies of $H_{k-1}$ while $H_{1}$ contains $2$ vertices. Thus, there are at least $4^{p}$ vertices in $G:=H_{p}$, which has treewidth $p$.
@@ 6 Complexity of Fixed-Ordering PQ-Layouts
We “cut” the circle at some point, which gives an interval representation where some vertices $S$ of $H$ refer to two intervals; in Figure 6, the circular arcs $a$, $b$, and $d$ are cut into intervals $a_{1}$ and $a_{2}$, $b_{1}$ and $b_{2}$, and $d_{1}$ and $d_{2}$, respectively. We let the endpoints of all intervals be the vertices of $G$, whose ordering $\sigma$ along the spine is taken from the ordering of the corresponding endpoints of the intervals (for the intervals belonging to $S$, the order is arbitrary). Further, for each interval, we have an edge in $G$ connecting its two endpoints. To obtain $w$, we assign to the edges of $G$ the numbers 1 to $m$, where $m=|V(H)|+|S|$, in decreasing order of the right endpoints of the intervals. This way, every pair of edges whose corresponding intervals share a common point need to be assigned to distinct priority queues.
@@
So far, there is no mechanism that assures that $a_{1}$ and $a_{2}$, $b_{1}$ and $b_{2}$, etc., are assigned to the same priority queue. To this end, we add, for each such pair $\langle a_{1},a_{2}\rangle$, an edge from the left endpoint of $a_{1}$ to the right endpoint of $a_{2}$. We call these $|S|$ new edges synchronization edges, and we assign them the weights $m+1,\dots,m+|S|$ in decreasing order of their right endpoints. Furthermore, we add $k-1,k-2,\dots$ heavy edges below the leftmost, second leftmost, … vertex on the spine, respectively, as well as, $k-1,k-2,\dots$ heavy edges below the rightmost, second rightmost, … vertex on the spine, respectively, see Figure 6. Clearly, $k\geq|S|$, otherwise we have a no-instance. The heavy edges have distinct right endpoints and their weights are chosen such that (i) they are greater than $M$, where $M=m+|S|$ is the largest weight used so far, and (ii) they are chosen in decreasing order from left to right such that each pair of heavy edges needs to be assigned to distinct priority queues if they overlap.
@@
If we have a PQ-layout of $\langle G,w\rangle$ under vertex ordering $\sigma$ with $k$ priority queues, we obtain, due to the choice of $w$ and $\sigma$, a coloring of $H$ by using the priority queues as colors. In particular, due to 2, both parts of the “cut” circular arcs get the same color. Conversely, we can easily assign the edges of $\langle G,w\rangle$ to $k$ priority queues if we are given a $k$-coloring of $H$. Clearly, our reduction can be implemented to run in polynomial time.
@@ 7 Open Problems
We conclude by summarizing a few intriguing problems that remain open. (i) Improve the upper and lower bounds for $\operatorname{pqn}(K_{n})$. (ii) Determine the complexity of deciding $\operatorname{pqn}(G)\leq k$ for $k>1$, or deciding $\operatorname{pqn}(G,w)\leq k$ for a given edge weighting $w$ when $k$ is a constant or no vertex ordering is given. (iii) Find graph families with bounded priority queue number; what about outerplanar graphs? (iv) Can the edge density of a graph $G$ be upper bounded by a term in $\operatorname{pqn}(G)$? (v) One can also investigate the structure of edge-weighted graphs $\langle G,w\rangle$ with a fixed vertex-ordering $\prec$. Edge-partitions of $G$ into $k$ priority queues are equivalent to proper $k$-colorings of an associated graph $H$, while $t$-inversions correspond to $t$-cliques in $H$. Is $\chi(H)$ bounded by a function in terms of $\omega(H)$? Is determining $\omega(H)$NP-complete?
@@
Appendix A Omitted Proofs of Section 3
@@
Let $\Gamma$ be any PQ-layout of $K_{n,n}$. We first show that there exists a subgraph $K_{\frac{n}{2},\frac{n}{2}}$ with independent sets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ of size $\frac{n}{2}$ such that the restriction $\Gamma^{\prime}$ of $\Gamma$ to the subgraph induced by $A^{\prime}\cup B^{\prime}$ is a separated PQ-layout. [^4] For simplicity, we assume that $n$ is even. Let $v_{1},\ldots,v_{t}$ be the first $t$ vertices in the linear ordering of $\Gamma$, such that |$|\{v_{1},\ldots,v_{t}\}\cap A|=\frac{n}{2}$ and $t$ is minimum. If |$|\{v_{t+1},\ldots,v_{2n}\}\cap B|\geq\frac{n}{2}$, we set $A^{\prime}=\{v_{1},\ldots,v_{t}\}\cap A$ and $B^{\prime}=\{v_{q},\ldots,v_{2n}\}\cap B$ where $q\geq t+1$ is chosen so that $|B^{\prime}|=\frac{n}{2}$. Otherwise, by the choice of $t$, we have that |$|\{v_{t+1},\ldots,v_{2n}\}\cap A|=\frac{n}{2}$, and, by the pigeon-hole principle, also |$|\{v_{1},\ldots,v_{t}\}\cap B|>\frac{n}{2}$. In this case, we choose $B^{\prime}=\{v_{1},\ldots,v_{p}\}\cap B$ where $p<t$ is chosen so that $|B^{\prime}|=\frac{n}{2}$ and $A^{\prime}=\{v_{t+1},\ldots,v_{2n}\}\cap A$. This proves the existence of $\Gamma^{\prime}$.
@@
Figure 7: Illustrations for the proof Theorem 3.3. In this matrix of size $n\times\frac{n}{2}$, all possible edge weights appear in the rows and the indices of the vertices $b_{1},\dots,b_{\frac{n}{2}}$ appear in the columns.
@@
We now show that for some $k$ there is a $k$-inversion, that is, there is a sequence of edges $e_{1},\ldots,e_{k}$ such that for $1\leq i<j\leq k$, $w(i)>w(j)$ and $b_{\eta(i)}<b_{\eta(j)}$ where $\eta(i)$ is the index of the endpoint of edge $e_{i}$ in $B^{\prime}$. Another way to see this is by visualizing the relation between $\eta(i)$ and $w(e_{i})$ as a binary matrix [^5] A similar, although not identical, representation is used by Alam et al. [35] to prove the mixed page number of $K_{n,n}$. $M$ of size $n\times\frac{n}{2}$ where $M[w,\eta]=1$ if there is an edge $e_{i}$ in $\Gamma^{\prime}$ with $w(e_{i})=w$ and $\eta(i)=\eta$, and $M[w,\eta]=0$ otherwise; see Figure 7. In this view of the problem, for a given edge $e_{i}$ (see the red cell in Figure 7(a)), forbidden configurations are created with any edge whose cell (with value $1$) occurs in the region $\mathcal{R}(e_{i})$ to the top-left of the cell of $e_{i}$ (see the red shaded cells in Figure 7(a)). Moreover, each column of the matrix has exactly $\frac{n}{2}$ cells with value $1$ (corresponding to the $\frac{n}{2}$ edges incident to each vertex in $B^{\prime}$). Our desired sequence $e_{1},\ldots,e_{k}$ thus is a strictly monotonically decreasing path through the matrix along cells with value $1$; see the red path in Figure 7(b). It is easy to see that each pair of edges along the sequence creates a forbidden configuration, i.e., $k$ priority queues are necessary.
@@
To compute a strictly monotonically decreasing path of maximum length, we proceed iteratively as follows. At step 1 we compute a subset of candidate edges (cells) for $e_{1}$, containing to every cell of an edge $e$ for which $\mathcal{R}(e)$ is empty (i.e., it does not contain cells of value 1). At step $i$, with $i\geq 2$, we compute a subset of candidate edges (cells) for $e_{i}$, containing every cells of an edge $e$ for which $\mathcal{R}(e)$ is empty after all cells of the candidate edges for $e_{j}$ with $j<i$ have been removed (i.e., are set to 0 in the matrix). For example, in Figure 7(c), the red cells correspond to the candidate edges for $e_{1}$, while the blue cells correspond to the candidate edges for $e_{2}$. We iteratively repeat the process until all $\frac{n^{2}}{4}$ cells with value $1$ have been set to 0. The number $k$ of steps in this process yields the length of the sequence and hence the required number of priority queues.
@@
We now compute the value of $k$. To this end, we first prove that, in step $i$, the topmost $i-1$ rows and the leftmost $i-1$ columns are empty (i.e., all cells are set to 0). The proof is by induction on $i$. When $i=1$ the claim is trivially true. Assume that the claim is true for $i-1$, with $i\geq 2$. For each edge $e$ whose cell is at column $i$, the region $\mathcal{R}(e)$ contains only cells at columns that are to the left of column $i$, which are empty by the inductive hypothesis. Hence, all cells at column $i$ are set to 0 in step $i$. An analogous argument applies for the topmost $i-1$ rows. Moreover, denote by $M^{\prime}$ the sub-matrix of $M$ obtained from $m$ after removing the $i-1$ leftmost columns and the $i-1$ topmost rows. In each diagonal of $M^{\prime}$ of slope $-1$, at most one non-empty cell has an empty region to the top-left. Observe that $M^{\prime}$ has $n-i+1$ rows and $\frac{n}{2}-i+1$ columns; thus the number of such diagonals is $\frac{3n}{2}-2i+1$ and the number of cells set to 0 at step $i$ is at most $\frac{3n}{2}-2i+1$. The number of iterations $k$ is the minimum integer that satisfies the following inequality:
@@
Appendix B Omitted Proofs of Section 4
See 4.1

### v3

@@ 4.3 Characterization and Recognition
Proof 4.19 (Proof Sketch).
@@
We analyze the structure of $G$ and we see that either $G$ is a graph that has $\operatorname{pqn}(G)=1$ due to a result from Section˜4.1, or $G$ has some $\mathcal{F}_{i}$ for $i\in\{1,\ldots,8\}$ as a minor.
@@
If $G$ is a tree, then $\operatorname{pqn}(G)=1$ by Lemma˜4.1. If $G$ contains exactly one cycle $C$, then GE(C)G-E(C) is a forest $F$, and we consider each component of $F$ as rooted at $C$. If every component is an isolated vertex or a rooted star, then $\operatorname{pqn}(G)=1$ by Lemma˜4.7. If some component is more complex than a rooted star or a rooted caterpillar, then $G$ contains $\mathcal{F}_{5}$ as a minor. Now suppose some component $K$ is a rooted caterpillar. If all other components are isolated vertices, then $\operatorname{pqn}(G)=1$ by Lemma˜4.9. If there are two further non-trivial components, then $G$ contains $\mathcal{F}_{6}$ as a minor. So assume that there is exactly one further non-trivial component $K^{\prime}\neq K$. If the roots of $K$ and $K^{\prime}$ have distance at least $3$ in one direction along $C$, then $G$ contains $\mathcal{F}_{4}$ as a minor. If none of the above applies, either $C$ is a triangle and $\operatorname{pqn}(G)=1$ by Lemma˜4.11, or $C$ is a quadrangle and the roots of $K$ and $K^{\prime}$ are opposite on $C$, and $\operatorname{pqn}(G)=1$ by Lemma˜4.11.
The case where $G$ contains at least two cycles remains. If $G$ has two edge-disjoint cycles, then $G$ contains $\mathcal{F}_{2}$ as a minor. If two cycles have at least two vertices but no edge in common, then $G$ contains $\mathcal{F}_{8}$ as a minor. If any two cycles share at least one edge, then either $G=K_{2,3}$ or $G=K_{4}-e$ and we have $\operatorname{pqn}(G)=1$ by Lemmas˜4.13 and 4.14, or $G$ contains $\mathcal{F}_{1}$, $\mathcal{F}_{3}$, $\mathcal{F}_{7}$, or $\mathcal{F}_{8}$ as a minor.
Corollary 4.20.
The family of graphs with priority queue number 1 is minor-closed.
@@
Theorem˜4.18 implies that the structure of a graph $G$ with $\operatorname{pqn}(G)=1$ is quite limited. In fact, forbidden minors $\mathcal{F}_{1},\mathcal{F}_{2}$ and $\mathcal{F}_{3}$ imply that $G$ can contain at most three cycles. More precisely, if $G$ contains two cycles, it is necessarily a $K_{2,3}$ or a $K_{4}-e$ since forbidden minors $\mathcal{F}_{7}$ and $\mathcal{F}_{8}$ forbid any other edge. Moreover, if $G$ contains a single cycle, forbidden minor $\mathcal{F}_{4}$ implies that no non-trivial tree (i.e., a tree that is more than a caterpillar) can be attached to the cycle. Finally, forbidden minors $\mathcal{F}_{4}$ and $\mathcal{F}_{6}$ dictate how legs can be combined with an attached caterpillar. $\mathcal{F}_{6}$ implies that at most one vertex of the cycle can have legs whereas $\mathcal{F}_{4}$ implies that, for a cycle of length $4$, legs can only be attached to the vertex opposite of the vertex attached to the caterpillar whereas for cycles of lengths larger than $4$, no legs are allowed. These observations in conjunction with Lemmas˜4.1, 4.3, 4.7, 4.11, 4.13 and 4.14 imply the following:
@@ 5 PQ-Layouts of Graphs with Bounded Pathwidth and Treewidth
We present results for graphs of bounded pathwidth and bounded treewidth. We assume familiarity with these concepts; otherwise see Appendix˜C for formal definitions.
@@
To define the partition of $E$ into $p+1$ priority queues $\mathcal{P}_{1},\ldots,\mathcal{P}_{p+1}$, we consider a proper vertex coloring of $G$ with $\chi(G)=\omega(G)=p+1$ colors. This exists, as $G$ is an interval graph. Now for $c=1,\ldots,p+1$ let $\mathcal{P}_{c}$ be the set of all edges in $G$ whose right endpoint in the vertex ordering has color $c$. In fact, each $\mathcal{P}_{c}$ contains none of the forbidden configurations in Figure˜2, and hence is a priority queue independent of the edge-weights, since in each forbidden configuration in Figure˜2 there would be an edge in $G$ between the right endpoints $v_{r}$ and $v_{r^{\prime}}$ of the two edges $e$ and $e^{\prime}$ forming the forbidden configuration. But this would imply different colors for $v_{r}$ and $v_{r^{\prime}}$ and therefore different priority queues for $e$ and $e^{\prime}$.
We remark that Corollary˜3.5 implies that there are graphs with pathwidth $p$ and priority queue number at least $\left\lfloor\frac{3-\sqrt{5}}{8}(p+1)\right\rfloor$ as $K_{p+1}$ has pathwidth $p$.
@@
We define a vertex ordering $\sigma$ of a $2$-tree rooted at $e^{*}=v^{\prime}v^{\prime\prime}$ to be left-growing if no vertex (different from $^{\prime}, v′′v^{\prime\prime}$) appears to the right of both its parents. First we show that it is enough to force a $p$-inversion (and hence $\operatorname{pqn}(G)\geq p$) in every left-growing vertex ordering.
@@
Figure 5: Illustration of the 22-tree $H_{k}$ with edge-weighting $w_{k}$, for $k\geq 2$.
We then construct a sequence $\langle H_{1},w_{1}\rangle,\langle H_{2},w_{2}\rangle,\ldots$ of edge-weighted rooted $2$-trees, so that for every $k$, every left-growing vertex ordering of $\langle H_{k},w_{k}\rangle$ contains a $k$-inversion. The construction of $\langle H_{k},w_{k}\rangle$ is recursive, starting with $\langle H_{1},w_{1}\rangle$ being just a single edge of any weight. For $k\geq 2$, we start with $d=k^{2}$ vertices $a_{1},\ldots,a_{d}$ stacked onto the root edge $e^{*}=v^{\prime}v^{\prime\prime}$, and stacking a vertex $b^{\prime}_{i}$ onto each $v^{\prime}a_{i}$, as well as a vertex $b^{\prime\prime}_{i}$ onto $v^{\prime\prime}a_{i}$ for $i=1,\ldots,d$. All these edges have weight 0. Then, each $b^{\prime}_{i}a_{i}$ and $b^{\prime\prime}_{i}a_{i}$ is used as the base edge of two separate copies of $\langle H_{k-1},w_{k-1}\rangle$, where however the edge-weights $w_{k-1}$ of each copy are scaled and shifted to be in a specific half-open interval $x,y)\subset\mathbb{R}$ depending on the current base edge ($b^{\prime}_{i}a_{i}$ or $b^{\prime\prime}_{i}a_{i}$). See Figure˜5 for an illustration.
By induction, there is a (k1)(k-1)-inversion in each such copy of $H_{k-1}$. Using the pigeon-hole principle, and the Erdős–Szekeres theorem, we then identify an edge $a_{i}v^{\prime}$ (or $a_{i}v^{\prime\prime}$) which forms a $k$-inversion either together with (k1)(k-1)-inversion of one copy of $H_{k-1}$, or with k1k-1 edges, each from the (k1)(k-1)-inversion of a different copy of $H_{k-1}$.
To finish the proof, it is then enough to take $\langle H_{p},w_{p}\rangle$ as constructed above, and then apply ˜1 to obtain the desired edge-weighted $2$-tree $\langle G_{p},w_{p}\rangle$ with $\operatorname{pqn}(G_{p},w_{p})\geq p$.
@@
We observe that $H_{k}$ in the proof of Theorem˜5.3 contains $4$ copies of $H_{k-1}$ while $H_{1}$ contains $2$ vertices. Thus, there are at least $4^{p}$ vertices in $G:=H_{p}$, which has treewidth $p$.
@@ 6 Complexity of Fixed-Ordering PQ-Layouts
We “cut” the circle at some point, which gives an interval representation where some vertices $S$ of $H$ refer to two intervals; in Figure˜6, the circular arcs $a$, $b$, and $d$ are cut into intervals $a_{1}$ and $a_{2}$, $b_{1}$ and $b_{2}$, and $d_{1}$ and $d_{2}$, respectively. We let the endpoints of all intervals be the vertices of $G$, whose ordering $\sigma$ along the spine is taken from the ordering of the corresponding endpoints of the intervals (for the intervals belonging to $S$, the order is arbitrary). Further, for each interval, we have an edge in $G$ connecting its two endpoints. To obtain $w$, we assign to the edges of $G$ the numbers 1 to $m$, where $m=|V(H)|+|S|$, in decreasing order of the right endpoints of the intervals. This way, every pair of edges whose corresponding intervals share a common point need to be assigned to distinct priority queues.
@@
So far, there is no mechanism that assures that $a_{1}$ and $a_{2}$, $b_{1}$ and $b_{2}$, etc., are assigned to the same priority queue. To this end, we add, for each such pair $\langle a_{1},a_{2}\rangle$, an edge from the left endpoint of $a_{1}$ to the right endpoint of $a_{2}$. We call these $|S|$ new edges synchronization edges, and we assign them the weights $m+1,\dots,m+|S|$ in decreasing order of their right endpoints. Furthermore, we add $k-1,k-2,\dots$ heavy edges below the leftmost, second leftmost, … vertex on the spine, respectively, as well as, $k-1,k-2,\dots$ heavy edges below the rightmost, second rightmost, … vertex on the spine, respectively, see Figure˜6. Clearly, $k\geq|S|$, otherwise we have a no-instance. The heavy edges have distinct right endpoints and their weights are chosen such that (i) they are greater than $M$, where $M=m+|S|$ is the largest weight used so far, and (ii) they are chosen in decreasing order from left to right such that each pair of heavy edges needs to be assigned to distinct priority queues if they overlap.
@@
If we have a PQ-layout of $\langle G,w\rangle$ under vertex ordering $\sigma$ with $k$ priority queues, we obtain, due to the choice of $w$ and $\sigma$, a coloring of $H$ by using the priority queues as colors. In particular, due to ˜2, both parts of the “cut” circular arcs get the same color. Conversely, we can easily assign the edges of $\langle G,w\rangle$ to $k$ priority queues if we are given a $k$-coloring of $H$. Clearly, our reduction can be implemented to run in polynomial time.
@@ 7 Open Problems
[…]
@@
Appendix A Omitted Proofs of Section˜3
@@
Let $\Gamma$ be any PQ-layout of $K_{n,n}$. We first show that there exists a subgraph $K_{\frac{n}{2},\frac{n}{2}}$ with independent sets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ of size $\frac{n}{2}$ such that the restriction $\Gamma^{\prime}$ of $\Gamma$ to the subgraph induced by $A^{\prime}\cup B^{\prime}$ is a separated PQ-layout. [^4] For simplicity, we assume that $n$ is even. Let $v_{1},\ldots,v_{t}$ be the first $t$ vertices in the linear ordering of $\Gamma$, such that $|\{v_{1},\ldots,v_{t}\}\cap A|=\frac{n}{2}$ and $t$ is minimum. If $|\{v_{t+1},\ldots,v_{2n}\}\cap B|\geq\frac{n}{2}$, we set $A^{\prime}=\{v_{1},\ldots,v_{t}\}\cap A$ and $B^{\prime}=\{v_{q},\ldots,v_{2n}\}\cap B$ where $q\geq t+1$ is chosen so that $|B^{\prime}|=\frac{n}{2}$. Otherwise, by the choice of $t$, we have that $|\{v_{t+1},\ldots,v_{2n}\}\cap A|=\frac{n}{2}$, and, by the pigeon-hole principle, also $|\{v_{1},\ldots,v_{t}\}\cap B|>\frac{n}{2}$. In this case, we choose $B^{\prime}=\{v_{1},\ldots,v_{p}\}\cap B$ where $p<t$ is chosen so that $|B^{\prime}|=\frac{n}{2}$ and $A^{\prime}=\{v_{t+1},\ldots,v_{2n}\}\cap A$. This proves the existence of $\Gamma^{\prime}$.
@@
Figure 7: Illustrations for the proof Theorem˜$3.$ In this matrix of size $n\times\frac{n}{2}$, all possible edge weights appear in the rows and the indices of the vertices $b_{1},\dots,b_{\frac{n}{2}}$ appear in the columns.
@@
We now show that for some $k$ there is a $k$-inversion, that is, there is a sequence of edges $e_{1},\ldots,e_{k}$ such that for $1\leq i<j\leq k$, $w(i)>w(j)$ and $b_{\eta(i)}<b_{\eta(j)}$ where $\eta(i)$ is the index of the endpoint of edge $e_{i}$ in $B^{\prime}$. Another way to see this is by visualizing the relation between $\eta(i)$ and $w(e_{i})$ as a binary matrix [^5] A similar, although not identical, representation is used by Alam et al. [35] to prove the mixed page number of $K_{n,n}$. $M$ of size $n\times\frac{n}{2}$ where $M[w,\eta]=1$ if there is an edge $e_{i}$ in $\Gamma^{\prime}$ with $w(e_{i})=w$ and $\eta(i)=\eta$, and $M[w,\eta]=0$ otherwise; see Figure˜7. In this view of the problem, for a given edge $e_{i}$ (see the red cell in Figure˜7(a)), forbidden configurations are created with any edge whose cell (with value $1$) occurs in the region $\mathcal{R}(e_{i})$ to the top-left of the cell of $e_{i}$ (see the red shaded cells in Figure˜7(a)). Moreover, each column of the matrix has exactly $\frac{n}{2}$ cells with value $1$ (corresponding to the $\frac{n}{2}$ edges incident to each vertex in $B^{\prime}$). Our desired sequence $e_{1},\ldots,e_{k}$ thus is a strictly monotonically decreasing path through the matrix along cells with value $1$; see the red path in Figure˜7(b). It is easy to see that each pair of edges along the sequence creates a forbidden configuration, i.e., $k$ priority queues are necessary.
@@
To compute a strictly monotonically decreasing path of maximum length, we proceed iteratively as follows. At step 1 we compute a subset of candidate edges (cells) for $e_{1}$, containing to every cell of an edge $e$ for which $\mathcal{R}(e)$ is empty (i.e., it does not contain cells of value 1). At step $i$, with $i\geq 2$, we compute a subset of candidate edges (cells) for $e_{i}$, containing every cells of an edge $e$ for which $\mathcal{R}(e)$ is empty after all cells of the candidate edges for $e_{j}$ with $j<i$ have been removed (i.e., are set to 0 in the matrix). For example, in Figure˜7(c), the red cells correspond to the candidate edges for $e_{1}$, while the blue cells correspond to the candidate edges for $e_{2}$. We iteratively repeat the process until all $\frac{n^{2}}{4}$ cells with value $1$ have been set to 0. The number $k$ of steps in this process yields the length of the sequence and hence the required number of priority queues.
@@
We now compute the value of $k$. To this end, we first prove that, in step $i$, the topmost i1i-1 rows and the leftmost i1i-1 columns are empty (i.e., all cells are set to 0). The proof is by induction on $i$. When $i=1$ the claim is trivially true. Assume that the claim is true for i1i-1, with $i\geq 2$. For each edge $e$ whose cell is at column $i$, the region $\mathcal{R}(e)$ contains only cells at columns that are to the left of column $i$, which are empty by the inductive hypothesis. Hence, all cells at column $i$ are set to 0 in step $i$. An analogous argument applies for the topmost i1i-1 rows. Moreover, denote by $M^{\prime}$ the sub-matrix of $M$ obtained from $m$ after removing the i1i-1 leftmost columns and the i1i-1 topmost rows. In each diagonal of $M^{\prime}$ of slope 1-1, at most one non-empty cell has an empty region to the top-left. Observe that $M^{\prime}$ has ni+1n-i+1 rows and $\frac{n}{2}-i+1$ columns; thus the number of such diagonals is $\frac{3n}{2}-2i+1$ and the number of cells set to 0 at step $i$ is at most $\frac{3n}{2}-2i+1$. The number of iterations $k$ is the minimum integer that satisfies the following inequality:
@@
Appendix B Omitted Proofs of Section˜4
See 4.7